paul is 3 years older than dave. daveis 4 years less than twice mike's age. how old are the three boysif their ages add up to 35
Answer by rfer(16322)

paul 2x-4+3
dave=2x-4
mike=x
x+2x-4+2x-4+3=35
5x-5=35
5x=35+5
x=40/5
x=8
2x-4=12
2x-4+3=15